OpenStudy (anonymous):

Solve the system of equations: 4x + 3y = 1 3x + 2y = 2

OpenStudy (anonymous):

1.multiply the 1st equation by 2 and 2nd equation by 3 2.subtract the 2nd equation from 1st equation and you will get 'x' value 3.substitute the 'x' value in one of the equation to get 'y' value

OpenStudy (anonymous):

4x+3y=1.....(1) 3x+2y=2....(2) (1)x2 8x+6y=2...(1)' (2)x3 9x+6y=6....(2)' (1)'-(2)' -x=-4 x=4 sub to (1) 4(4)+3y=1 16+3y=1 3y=15 y=5 (4,5)